OpenStudy (anonymous):

If 62.7 g N2 react with 23.8 g H2 according to the reaction below, how many grams of ammonia (NH3) can be produced, and how many grams of the excess reactant will be left over? Unbalanced equation: N2 + H2 → NH3 Be sure to show all of your work.

OpenStudy (anonymous):

First thing to do is to balance the equation, then maybe convert masses in moles.

OpenStudy (anonymous):

N2 + 3H2 ----> 2NH3 and 62.7/ 28 =2.24 and 23.8/2.02

OpenStudy (anonymous):

Idk what to do after tht

OpenStudy (anonymous):

which one is limiting?

OpenStudy (anonymous):

N2 is limiting right?

OpenStudy (anonymous):

yep so if you have one mole of nitrogen for every 2 moles of ammonia how many moles of ammonia do you have?

OpenStudy (anonymous):

I am not sure! I am confused!?

OpenStudy (anonymous):

The rest is calculation, 1 mole of N2 can react with 3 moles of H2 to make 2 moles of NH3

OpenStudy (anonymous):

can you write it on in numbers so it is easier to understand?

OpenStudy (anonymous):

in the equation. it says N2 +3H3 goes to 2NH3. you have the moles of N2 so use the ratio.\[2.24moles \ N_2 \times {2moles NH_3 \over 1mole N_2} = moles \ of \ ammonia.\]

OpenStudy (anonymous):

You have to get used to the thinking, it's like a cake recipe, except it's not eggs but moles !

OpenStudy (anonymous):

4.48 moles N2 produced

OpenStudy (anonymous):

Now since you have moles of ammonia use it's molar mass to find the grams. You found 4.48 moles NH3 not N2

OpenStudy (anonymous):

how do you convert 4.48 moles of NH3

OpenStudy (anonymous):

The molar mass of ammonia is 17.04grams per mole. found by adding the molar mass of each one so three H's and one N. Then use our molar mass equation.\[Molar Mass = {grams \over moles} \]therefore \[grams = moles \times molar\ mass\]\[grams\ of \ ammonia = 4.48 mol \times 17.04 {g \over mol}\]see how the units cancel to give grams?

OpenStudy (anonymous):

yes so 4.48 * 17.04 = 76.3392

OpenStudy (anonymous):

right. Now how many grams did we have in the first place?

OpenStudy (anonymous):

62.7 grams of N2 and 23.8 g H2

OpenStudy (anonymous):

add them together and you get the total grams. Subtract that from the grams of ammonia you found earlier and you'll get the excess hydrogen in grams. We know it's hydrogen becuase you used up all the nitrogen earlier because it was limiting.

OpenStudy (anonymous):

oh ok so its 86.5- 76.29= 10.21

OpenStudy (anonymous):

Thank you very much for all the help!!!

OpenStudy (anonymous):

no problem.